# A Scalable Diagonalization Framework for Tensor-Product Bitstring Selected Configuration Interaction

### 05 Mar 2026 v4

@@ I Introduction
Starting from Eq. (1), one of the most direct approximations is to retain only the important determinants—those carrying the large weights ($c_{K}^{2}$) in the wavefunction—an idea that underlies selected configuration interaction (SCI) methods. Over the past decades, several SCI methods have been proposed and developed, including configuration interaction using a perturbative selection made iteratively (CIPSI), [8] sparse FCI, [9] adaptive sampling CI (ASCI), [10, 11] semistochastic heat-bath CI (SHCI), [12, 13, 14] and iterative CI (iCI). [15, 16, 17] These approaches have demonstrated that such selection strategies can yield highly compact wavefunctions for a wide range of molecular systems, enabling near-FCI accuracy with only a small fraction of the FCI determinants.
From an algorithmic perspective, a generic SCI calculation consists of three stages: (1) selecting important determinants, (2) diagonalizing the CI Hamiltonian within the selected subspace, and (3) applying perturbative corrections to recover residual correlation. Leveraging modern supercomputing architectures, several SCI methods [8, 18, 10, 11, 12, 13, 14, 15, 16, 17] have implemented distributed-memory parallelization across all three stages, enabling determinant spaces to grow substantially beyond what was previously feasible. Yet in most existing implementations the CI vector remains fully replicated across nodes—not because the SCI framework inherently precludes distributed CI-vector storage, but because maintaining both efficiency and scalability with distributed CI-vector storage at large scale remains a nontrivial algorithmic and implementation challenge. As a consequence, reported SCI calculations typically involve up to $2\times 10^{9}$ determinants in stage (2). As a variational method, SCI systematically improves the wavefunction description as more determinants are retained. For modern SCI calculations targeting increasingly accurate descriptions of large-scale strongly correlated systems, the distributed diagonalization of extremely large selected determinant subspaces therefore becomes the central scalability bottleneck.
@@ II.3 Efficient Hamiltonian computation tailored for TBSCI
Although only a subset of determinants in the TPB determinant space $\mathcal{D}_{\mathrm{TPB}}$ is retained, TBSCI preserves the tensor-product indexing structure associated with the TPB representation. Importantly, the TPB structure serves as an organizing principle for indexing and Hamiltonian traversal; it does not require the retained determinants to form a tensor-product closure of $\mathcal{D}_{\mathrm{TPB}}$. As a consequence, for a fixed pair $(S_{w}^{\alpha},S_{x}^{\alpha})$, all $\beta$-side excitation candidates can be generated within the selected $\beta$-bitstring set $\{S^{\beta}\}$ and then filtered by the corresponding $\beta$-index sets $\mathcal{I}_{w}^{(\beta)}$ and $\mathcal{I}_{x}^{(\beta)}$. This property allows excitation relations among $\beta$-bitstrings to be precomputed once and reused across all $\alpha$-string pairs under the TPB indexing structure. Specifically, we construct the BETA_SINGLE_LINK and BETA_DOUBLE_LINK link tables (implemented as arrays), which store single and double excitation connectivity within $\{S^{\beta}\}$. As a result, for each admissible $(S_{w}^{\alpha},S_{x}^{\alpha})$ pair, the allowed Hamiltonian contributions are evaluated by traversing excitation links within $\{S^{\beta}\}$ rather than enumerating determinant pairs in the full product space.
@@ 28                    $S_{u}^{\beta}\leftarrow S_{v}^{\beta}$
The effective sparsity of single and double excitations can be approximately estimated as $\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, where $N_{\mathrm{TBSCI}}$ is the number of determinants retained in TBSCI, reflecting the reduction in the number of retained $\beta$-bitstrings relative to the FCI space. Accordingly, the memory usage of BETA_SINGLE_LINK and BETA_DOUBLE_LINK can be estimated to scale as $l_{\beta}N_{\mathrm{occ}}N_{\mathrm{vir}}\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$ and $l_{\beta}N_{\mathrm{occ}}^{2}N_{\mathrm{vir}}^{2}\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, respectively, for $N_{\mathrm{occ}}$ occupied and $N_{\mathrm{vir}}$ virtual orbitals. The BETA_DOUBLE_LINK array can be very large and is therefore partitioned into batches, one batch per process, with each batch containing a subset of $\{S_{u}^{\beta}\mid u=1,2,\dots,l_{\beta}\}$. For each $S_{u}^{\beta}$, its column BETA_DOUBLE_LINK(:, $S_{u}^{\beta}$) is precomputed and stored entirely on a single process with auxiliary arrays recording the corresponding index ranges to enable fast access. The computation of the [0,2] term is thus carried out batch by batch during one Davidson iteration (Lines 27–32 of Algorithm 1). Since this [0,2] term involves only a loop over $S_{w}^{\alpha}$ and does not reference $S_{x}^{\alpha}$, each process fetches the relevant BETA_DOUBLE_LINK batch only once per step, removing memory pressure while introducing negligible communication overhead.
@@
In Algorithm 1, the sparsity factor ($\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$) appears once in the computations of the $[2,0]$ and $[0,2]$ terms, but twice in that of the $[1,1]$ term. Consequently, the dominant contributions arise from the $[2,0]$ and $[0,2]$ terms, each scaling as $N_{\mathrm{TBSCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, whereas the $[1,1]$ term scales as $N_{\mathrm{TBSCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot(N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}})$. In this work, the TBSCI program is also employed to perform FCI calculations, where the sparsity factor becomes exactly one and the computational dependence simplifies to $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, with dominant contributions arising from the $[2,0]$, $[1,1]$, and $[0,2]$ terms. A summary of the scaling behavior and the dominant terms is provided in Table 1. It should be noted that the traditional Handy–Knowles determinant-based FCI algorithm scales as $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$[1, 2] which is more favorable for complete active-space calculations where the full determinant space is retained. More recently, the small-tensor-product distributed active space (STP-DAS) framework [7] has further extended FCI calculations to the quadrillion-determinant regime by introducing distributed active-space decompositions together with lossless compression of CI vectors, while maintaining high numerical accuracy. The present work is not intended to compete with existing FCI methodologies, [1, 2, 4, 6, 7] but to employ the FCI workload as a communication-intensive stress test for the distributed TBSCI eigensolver. As already reflected in Algorithm 1, the method is designed to operate on arbitrary determinant subsets of $\mathcal{D}_{\mathrm{TPB}}$. Its objective is therefore fundamentally different: rather than accelerating FCI, it enables scalable diagonalization of extremely large selected determinant spaces under fully distributed CI-vector storage.
@@
In existing SCI implementations that evaluate Hamiltonian elements on-the-fly, the mixed-spin $[1,1]$ term typically dominates the computational cost. For a fixed $\alpha$-pair with $\mathrm{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=1$, evaluating the $[1,1]$ term requires looping over all pairs $(S_{u}^{\beta},S_{v}^{\beta})$ with $S_{u}^{\beta}\in\mathcal{I}_{w}^{(\beta)}$ and $S_{v}^{\beta}\in\mathcal{I}_{x}^{(\beta)}$, which leads to a quadratic traversal over the $\beta$-bitstring space. In contrast, by traversing precomputed excitation connectivity within the selected $\beta$-bitstring set $\{S^{\beta}\}$, the dominant computational cost grows approximately with the number of retained determinants (as shown in Table 1), because determinant pairs are no longer enumerated explicitly, and each retained determinant participates in only a limited number of excitation connections determined by the Slater–Condon excitation structure. We note that when $|\mathcal{I}_{w}^{(\beta)}|$ is small relative to the typical excitation connectivity size, the constant-factor overhead associated with connectivity traversal and membership testing may exceed that of direct pair enumeration. In the present implementation, however, each $\mathcal{I}_{w}^{(\beta)}$ forms a contiguous interval, so membership testing reduces to a simple interval check and the above constant-factor consideration does not arise.
@@ II.4 MPI Communication Optimization Strategies
As described above, the distributed matrix–vector multiplication involves frequent inter-node data transfers. Although the underlying SCI formulation is not inherently tied to a specific communication strategy, communication contention can dominate the execution time at large node counts, resulting in degraded MPI efficiency and, in extreme cases, increased wall time per iteration. To maintain parallel efficiency at large node counts, we developed and integrated a suite of MPI communication optimization strategies within the distributed TBSCI eigensolver.
@@
In addition to excitation-based pruning, molecular symmetry provides a further mechanism for eliminating unnecessary communication. Specifically, if $\mathrm{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=2$ and $S_{w}^{\alpha}$ and $S_{x}^{\alpha}$ belong to different irreducible representations, the corresponding Hamiltonian matrix elements vanish identically, and the associated CI vector segments do not need to be fetched. Enforcing D2h symmetry, for instance, reduces communication and computational costs by a factor of $\sim 64$.
@@
Algorithm 1 allows estimation of computational costs associated with each $\alpha$-bitstring segment. The total workload of a process is approximated by summing the costs of its assigned bitstrings. While $\alpha$-bitstrings are assigned in ascending excitation order, balancing memory usage and computational load simultaneously is challenging. We define two expansion factors: the memory expansion factor (the ratio of assigned bitstring count to the average) and the computation expansion factor (the ratio of estimated cost to the average). Strictly enforcing memory balance (memory expansion factor $\approx 1$) risks severe computational imbalance, causing idle processes, whereas strictly enforcing computational balance (computation expansion factor $\approx 1$) may overload memory on some nodes. We therefore adopt a compromise strategy, maintaining both factors moderately above one to balance memory and computation effectively. For FCI calculations, where per-segment costs are uniform, a uniform distribution naturally balances both aspects.
@@ III.1 Scalability and Parallel Efficiency
Figure 2 presents the wall time of a single distributed matrix–vector multiplication—the dominant cost in Davidson diagonalization—across increasing node counts for all systems. Detailed timings and node-hour costs are provided in Table S1 in the Supplementary Material. Due to the substantial computational cost of these large-scale runs, the timing metrics are reported from the initial production run rather than averaged over multiple trials. For N2 (aug-cc-pVDZ) and CN (cc-pVTZ), the maximum and average wall times remain very close across all node counts, closely following the ideal strong-scaling trend. This indicates that communication delays remain minor and that the distributed eigensolver operates in a computation-dominated regime. For Cr2 (STO-3G), communication delays become increasingly significant at higher node counts. Although the maximum wall time decreases when increasing the node count from 5760 to 11520, a further increase to 14400 nodes results in a noticeable rise in the maximum wall time, indicating that communication contention begins to limit further strong-scaling gains for this system. The largest test is N2 (cc-pVTZ), involving more than $2.6\times 10^{12}$ determinants. Although the gap between the maximum and average wall times gradually increases with node count, the overall wall time continues to decrease up to 54,000 nodes (more than 2.5 million cores on Fugaku). Even at the largest scale tested, communication overhead increases substantially, yet the iteration cost remains primarily determined by computational work. Interestingly, despite being the largest system in terms of determinant count, N2 (cc-pVTZ) does not exhibit the most severe relative communication delays, suggesting that runtime communication behavior depends not only on system size but also on detailed excitation structure and workload distribution.
@@
All reported timings correspond to a single Davidson iteration. In practice, once the diagonalization is launched, the per-iteration wall time remains nearly constant throughout the iterative procedure. Consequently, the total runtime can be reliably estimated from the per-iteration cost and the number of Davidson iterations required for convergence, indicating stable runtime characteristics of the distributed TBSCI implementation at large node counts. In the present work we employ the standard Davidson algorithm as described in Ref. 18, without additional acceleration techniques. As a representative example, along the PES of N2 (aug-cc-pVDZ), the number of Davidson iterations required for convergence increases significantly as the bond is stretched. At $R=2.068$ a.u., convergence is achieved in 12 iterations, with a per-iteration wall time of about 698 s on 1,440 nodes, corresponding to a total wall time of approximately 2 hours. At $R=4.2$ a.u., where convergence is slower, 52 Davidson iterations are required, with a per-iteration wall time of about 420 s on 2,400 nodes, giving a total wall time of approximately 6 hours.

### 29 Apr 2026 v5

@@ I Introduction
Starting from Eq. (1), one of the most direct approximations is to retain only the important determinants—those carrying the large weights ($c_{K}^{2}$) in the wavefunction—an idea that underlies selected configuration interaction (SCI) methods. Over the past decades, several SCI methods have been proposed and developed, including configuration interaction using a perturbative selection made iteratively (CIPSI), [8] sparse FCI, [9] adaptive CI (ACI), [10] adaptive sampling CI (ASCI), [11, 12] semistochastic heat-bath CI (SHCI), [13, 14, 15] and iterative CI (iCI). [16, 17, 18] Related ideas have also been explored in approaches based on a priori truncation, [19, 20] as well as in more recent machine learning-based determinant selection strategies. [21, 22, 23, 24] These approaches have demonstrated that such selection strategies can yield highly compact wavefunctions for a wide range of molecular systems, enabling near-FCI accuracy with only a small fraction of the FCI determinants.
From an algorithmic perspective, a generic SCI calculation consists of three stages: (1) selecting important determinants, (2) diagonalizing the CI Hamiltonian within the selected subspace, and (3) applying perturbative corrections to recover residual correlation. Leveraging modern supercomputing architectures, several SCI methods [8, 25, 11, 12, 13, 14, 15, 16, 17, 18] have implemented distributed-memory parallelization across all three stages, enabling determinant spaces to grow substantially beyond what was previously feasible. Yet in most existing implementations the CI vector remains fully replicated across nodes—not because the SCI framework inherently precludes distributed CI-vector storage, but because maintaining both efficiency and scalability with distributed CI-vector storage at large scale remains a nontrivial algorithmic and implementation challenge. As a consequence, reported SCI calculations typically involve up to $2\times 10^{9}$ determinants in stage (2). [15] As a variational method, SCI systematically improves the wavefunction description as more determinants are retained. For modern SCI calculations targeting increasingly accurate descriptions of large-scale strongly correlated systems, the distributed diagonalization of extremely large selected determinant subspaces therefore becomes the central scalability bottleneck.
@@ II.3 Efficient Hamiltonian computation tailored for TBSCI
Although only a subset of determinants in the TPB determinant space $\mathcal{D}_{\mathrm{TPB}}$ is retained, TBSCI preserves the tensor-product indexing structure associated with the TPB representation. Importantly, the TPB structure serves as an organizing principle for indexing and Hamiltonian traversal; it does not require the retained determinants to form a tensor-product closure of $\mathcal{D}_{\mathrm{TPB}}$. As a consequence, for a fixed pair $(S_{w}^{\alpha},S_{x}^{\alpha})$, all $\beta$-side excitation candidates can be generated within the selected $\beta$-bitstring set $\{S^{\beta}\}$ and then filtered by the corresponding $\beta$-index sets $\mathcal{I}_{w}^{(\beta)}$ and $\mathcal{I}_{x}^{(\beta)}$. This property allows excitation relations among $\beta$-bitstrings to be precomputed once and reused across all $\alpha$-bitstring pairs under the TPB indexing structure. Specifically, we construct the BETA_SINGLE_LINK and BETA_DOUBLE_LINK link tables (implemented as arrays), which store single and double excitation connectivity within $\{S^{\beta}\}$. As a result, for each admissible $(S_{w}^{\alpha},S_{x}^{\alpha})$ pair, the allowed Hamiltonian contributions are evaluated by traversing excitation links within $\{S^{\beta}\}$ rather than enumerating determinant pairs in the full product space.
@@ 28                    $S_{u}^{\beta}\leftarrow S_{v}^{\beta}$
The effective sparsity of single and double excitations can be approximately estimated as $\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, where $N_{\mathrm{TBSCI}}$ is the number of determinants retained in TBSCI. Accordingly, the memory usage of BETA_SINGLE_LINK and BETA_DOUBLE_LINK can be estimated to scale as $l_{\beta}N_{\mathrm{occ}}N_{\mathrm{vir}}\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$ and $l_{\beta}N_{\mathrm{occ}}^{2}N_{\mathrm{vir}}^{2}\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, respectively, for $N_{\mathrm{occ}}$ occupied and $N_{\mathrm{vir}}$ virtual orbitals. The BETA_DOUBLE_LINK array can be very large and is therefore partitioned into batches, one batch per process, with each batch containing a subset of $\{S^{\beta}\}$. For each $S_{u}^{\beta}$, its column BETA_DOUBLE_LINK(:, $S_{u}^{\beta}$) is precomputed and stored entirely on a single process with auxiliary arrays recording the corresponding index ranges to enable fast access. The computation of the [0,2] term is thus carried out batch by batch during one Davidson iteration (Lines 27–32 of Algorithm 1). Since this [0,2] term involves only a loop over $S_{w}^{\alpha}$ and does not reference $S_{x}^{\alpha}$, each process fetches the relevant BETA_DOUBLE_LINK batch only once per step, removing memory pressure while introducing negligible additional communication volume.
@@
In Algorithm 1, the sparsity factor ($\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$) appears once in the computations of the $[2,0]$ and $[0,2]$ terms, but twice in that of the $[1,1]$ term. Consequently, the dominant contributions arise from the $[2,0]$ and $[0,2]$ terms, each scaling as $N_{\mathrm{TBSCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}}}$, whereas the $[1,1]$ term scales as $N_{\mathrm{TBSCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot(N_{\mathrm{TBSCI}}/N_{\mathrm{FCI}})$. In this work, the TBSCI program is also employed to perform FCI calculations, where the sparsity factor becomes exactly one and the computational dependence simplifies to $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, with dominant contributions arising from the $[2,0]$, $[1,1]$, and $[0,2]$ terms. A summary of the scaling behavior and the dominant terms is provided in Table 1. It should be noted that the traditional Handy–Knowles determinant-based FCI algorithm scales as $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$[1, 2] which is more favorable for complete active-space calculations where the full determinant space is retained. More recently, the small-tensor-product distributed active space (STP-DAS) framework [7] has further extended FCI calculations to the quadrillion-determinant regime by introducing distributed active-space decompositions together with lossless compression of CI vectors, while maintaining high numerical accuracy. The present work is not intended to compete with existing FCI methodologies, [1, 2, 3, 4, 5, 6, 7] but to employ the FCI workload as a communication-intensive stress test for the distributed TBSCI eigensolver. As already reflected in Algorithm 1, the method is designed to operate on arbitrary determinant subsets of $\mathcal{D}_{\mathrm{TPB}}$. Its objective is therefore fundamentally different: rather than accelerating FCI, it enables scalable diagonalization of extremely large selected determinant spaces under fully distributed CI-vector storage.
@@
In existing SCI implementations that evaluate Hamiltonian elements on-the-fly, the mixed-spin $[1,1]$ term typically dominates the computational cost. For a fixed $\alpha$-pair with $\mathrm{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=1$, evaluating the $[1,1]$ term requires looping over all pairs $(S_{u}^{\beta},S_{v}^{\beta})$ with $u\in\mathcal{I}_{w}^{(\beta)}$ and $v\in\mathcal{I}_{x}^{(\beta)}$, which leads to a quadratic traversal over the corresponding $\beta$-bitstring index sets. In contrast, by traversing precomputed excitation connectivity within the selected $\beta$-bitstring set $\{S^{\beta}\}$, the dominant computational cost grows approximately in proportion to the number of retained determinants (as shown in Table 1), because determinant pairs are no longer enumerated explicitly and each retained determinant participates in only a limited number of excitation connections determined by the Slater–Condon excitation structure. We note that when $|\mathcal{I}_{w}^{(\beta)}|$ is small relative to the typical excitation connectivity size, the constant-factor overhead associated with connectivity traversal and membership testing may exceed that of direct pair enumeration. In the present implementation, however, each $\mathcal{I}_{w}^{(\beta)}$ forms a contiguous interval, so membership testing reduces to a simple interval check and the above constant-factor consideration does not arise.
@@ II.4 MPI Communication Optimization Strategies
As described above, the distributed matrix–vector multiplication involves frequent inter-node data transfers. Although the underlying SCI formulation is not inherently tied to a specific communication strategy, communication contention can dominate the execution time at large node counts, resulting in degraded MPI efficiency and, in extreme cases, even an increase in wall time with increasing node count. To maintain parallel efficiency at large node counts, we developed and integrated a suite of MPI communication optimization strategies within the distributed TBSCI eigensolver.
@@
In addition to excitation-based pruning, molecular symmetry provides a further mechanism for eliminating unnecessary communication. Specifically, if $\mathrm{DIFF}(S_{w}^{\alpha},S_{x}^{\alpha})=2$ and $S_{w}^{\alpha}$ and $S_{x}^{\alpha}$ belong to different irreducible representations, the corresponding Hamiltonian matrix elements vanish identically, and the associated CI vector segments do not need to be fetched. Exploiting D2h symmetry, for instance, reduces communication and computational costs by a factor of $\sim 64$.
@@
Algorithm 1 allows estimation of computational costs associated with each $\alpha$-bitstring segment. The total workload of a process is approximated by summing the costs of its assigned $\alpha$-bitstrings. While $\alpha$-bitstrings are assigned in ascending excitation order, balancing memory usage and computational load simultaneously is challenging. We define two expansion factors: the memory expansion factor (the ratio of assigned bitstring count to the average) and the computation expansion factor (the ratio of estimated cost to the average). Strictly enforcing memory balance (memory expansion factor $\approx 1$) risks severe computational imbalance, causing idle processes, whereas strictly enforcing computational balance (computation expansion factor $\approx 1$) may overload memory on some nodes. We therefore adopt a compromise strategy, maintaining both factors moderately above one to balance memory and computation effectively. For FCI calculations, where per-segment costs are uniform, a uniform distribution naturally balances both aspects.
@@ III.1 Scalability and Parallel Efficiency
Figure 2 presents the wall time of a single distributed matrix–vector multiplication—the dominant cost in Davidson diagonalization—across increasing node counts for all systems. Detailed timings and node-hour costs are provided in Table S1 in the Supplementary Material. Due to the substantial computational cost of these large-scale runs, the timing metrics are reported from the initial production run rather than averaged over multiple trials. For N2 (aug-cc-pVDZ) and CN (cc-pVTZ), the maximum and average wall times remain very close across all node counts, closely following the ideal strong-scaling trend. This indicates that communication delays remain minor and that the distributed eigensolver operates in a computation-dominated regime. For Cr2 (STO-3G), communication delays become increasingly significant at higher node counts. Although the maximum wall time decreases when increasing the node count from 5760 to 11520, a further increase to 14400 nodes results in a noticeable rise in the maximum wall time, indicating that communication contention begins to limit further strong-scaling gains for this system. The largest test is N2 (cc-pVTZ), involving more than $2.6\times 10^{12}$ determinants. Although the gap between the maximum and average wall times generally increases with node count, the overall wall time continues to decrease up to 54,000 nodes (more than 2.5 million cores on Fugaku). Even at the largest scale tested, communication overhead increases substantially, yet the iteration cost remains primarily determined by computational work. Interestingly, despite being the largest system in terms of determinant count, N2 (cc-pVTZ) does not exhibit the most severe relative communication delays, suggesting that runtime communication behavior depends not only on system size but also on detailed excitation structure and workload distribution.
@@
All reported timings correspond to a single Davidson iteration. In practice, once the diagonalization is launched, the per-iteration wall time remains nearly constant throughout the iterative procedure. Consequently, the total runtime can be reliably estimated from the per-iteration cost and the number of Davidson iterations required for convergence, indicating stable runtime characteristics of the distributed TBSCI implementation at large node counts. In the present work we employ the standard Davidson algorithm as described in Ref. 25, without additional acceleration techniques. As a representative example, along the PES of N2 (aug-cc-pVDZ), the number of Davidson iterations required for convergence increases significantly as the bond is stretched. At $R=2.068$ a.u., convergence is achieved in 12 iterations, with a per-iteration wall time of about 698 s on 1,440 nodes, corresponding to a total wall time of approximately 2 hours. At $R=4.2$ a.u., where convergence is slower, 52 Davidson iterations are required, with a per-iteration wall time of about 420 s on 2,400 nodes, giving a total wall time of approximately 6 hours.